# Apparent delay of the Kibble-Zurek mechanism in quenched open systems

### v1

@@ Abstract
We report a new intermediate regime in the quench time, $\tau_{q}$, separating the usual validity of the Kibble-Zurek mechanism (KZM) and its breakdown for rapid quenches in open systems under finite quench protocols. It manifests in the power-law scaling of the transition time with $\tau_{q}$ as the system appears to enter the adiabatic regime, even though the ramp is already terminated and the final quench value is held constant. This intermediate regime, which we dub as the delayed KZM, emerges due to the dissipation preventing the system from freezing in the impulse regime. This results in a large delay between the actual time the system undergoes a phase transition and the time inferred from a threshold-based criterion for the order parameter, as done in most experiments. We demonstrate using the open Dicke model and its one-dimensional lattice version that this phenomenon is a generic feature of open systems that can be mapped onto an effective coupled oscillator model. We also show that the phenomenon becomes more prominent near criticality, and its effects on the transition time measurement can be further exacerbated by large threshold values for an order parameter. Due to this, we propose an alternative method for threshold-based criterion which uses the spatio-temporal information, such as the system’s defect number, for identifying the transition time.
@@ I Introduction
Initially formulated to describe the evolution of topological defects in the early universe [1, 2, 3], the KZM has been successful in describing the dependence of the defect number and duration of a continuous phase transition on the quench timescale, $\tau_{q}$ [4]. In particular, the theory has been tested in multiple platforms, ranging from atomic Bose-Einstein condensates [5, 6, 7, 8, 9, 10, 11, 12], spin systems [13, 14, 15, 16], Rydberg atom set-ups [17, 18], and trapped-ion systems [19, 20]. It has also been tested in dissipative quantum systems [21, 22, 23, 24, 25, 26, 16, 27], and has recently been extended to include generic non-equilibrium systems [28, 29, 30].
Under the standard KZM, a generic closed system with a continuous phase transition has a diverging relaxation time, $\tau$, and correlation length, $\xi$, as it approaches its critical point, $\lambda_{c}$. In particular, one expects $\tau$ and $\xi$ to scale as $\tau\propto|\varepsilon|^{-vz}$ and $\xi\propto|\varepsilon|^{-v}$ [4], respectively, where $\varepsilon=\left(\lambda-\lambda_{c}\right)/\lambda_{c}$ is the reduced distance of the control parameter, $\lambda$, from the critical point, while $v$ and $z$ are the static and dynamic critical exponents, respectively. It is then expected that if the system is linearly quenched via a ramp protocol, $\varepsilon=t/\tau_{q}$, the system will become frozen near $\lambda_{c}$ due to $\tau$ diverging. This motivates the introduction of the Adiabatic-Impulse (AI) approximation, where the system’s dynamics are classified into two regimes [4]. Far from $\lambda_{c}$, the system is in an adiabatic regime, in which its macroscopic quantities adiabatically follow the quench. Near $\lambda_{c}$, the system enters the impulse regime, wherein all relevant observables remain frozen even after passing $\lambda_{c}$. It only re-enters the adiabatic regime and transitions to a new phase after some finite time referred to as the freeze-out time, $\hat{t}$, has passed [4]. This occurs after the system reaches the AI crossover point, $\varepsilon(\hat{t})$, setting $\hat{t}\sim\tau\left(\varepsilon(\hat{t})\right)$ [4]. The KZM predicts that, due to the scaling of $\tau$, $\hat{t}$ and $\varepsilon(\hat{t})$ must follow the scaling laws [4]
@@
Measuring the exact value of $\hat{t}$ and $\varepsilon(\hat{t})$ is a non-trivial task unlike defect counting due to the limitations in detecting the exact time a system re-enters the adiabatic regime. As such, it is common to employ a threshold criterion and measure instead the transition time, $\hat{t}_{\mathrm{th}}$, which is the time it takes for an order parameter to reach a given threshold after passing $\lambda_{c}$. The crossover point at the transition time, $\varepsilon(\hat{t}_{\mathrm{th}})$, is similarly defined. For a sufficiently small threshold value, it is assumed that $\hat{t}_{\mathrm{th}}$ is a good approximation for $\hat{t}$. While this method is successful in showing the power-law scaling of $\hat{t}_{\mathrm{th}}$ and $\varepsilon(\hat{t}_{\mathrm{th}})$ as a function of $\tau_{q}$ [5, 12, 36, 10, 31, 27, 40], it remains unclear whether the inherent deviation between $\hat{t}$ and $\hat{t}_{\mathrm{th}}$ does not lead to any significant effects on the scaling of the KZM quantities for any generic quench protocols.
In this work, we report an intermediate regime separating the breakdown and validity of the KZM appearing in open systems under a finite quench protocol depicted in Fig. 1(a). In this regime, the transition time follows the power-law scaling predicted by the KZM even though the system appears to relax after the quench has terminated, as illustrated in Fig. 1(b). As we will show later, this regime manifests precisely due to the dissipation exacerbating the deviation between the freeze-out time and the transition time, leading to a delay in the detection of the phase transition, as schematically represented in Fig. 1(b). We demonstrate using the open Dicke model (DM) [41, 42] and its one-dimensional lattice extension, the open Dicke lattice model (DLM) [43] that the range of $\tau_{q}$ where we observe this ”delayed” KZM is a generic feature of open systems with finite dissipation strength, $\kappa$. We also show that the delayed KZM is more prominent near criticality and that its signatures become more significant for large threshold values for an order parameter. Thus, our work highlights subtleties of the KZM in open systems in finite quench scenarios relevant to experiments.
@@ II Delayed KZM: Theory
where $\mathcal{D}\hat{\rho}=\sum_{\ell}\kappa_{\ell}\left(2\hat{L}_{\ell}\hat{\rho}\hat{L}_{\ell}^{\dagger}-\left\{\hat{L}^{\dagger}_{\ell}\hat{L}_{\ell},\hat{\rho}\right\}\right)$ is the dissipator, and $\hat{H}(\varepsilon(t))$ is the time-dependent Hamiltonian of the system. The system undergoes a phase transition from a normal phase (NP), in which the global symmetry of the system is preserved, to a symmetry-broken phase via a finite quench protocol,
$\varepsilon(t)=\begin{cases}t/\tau_{q}\;\;\;&t_{i}\leq t\leq\varepsilon_{f}\tau_{q}\\ \varepsilon_{f}\;\;\;&\varepsilon_{f}\tau_{q}<t\leq t_{f}\end{cases},$ (3)
@@
where $\omega$ and $\omega_{0}$ are the transition frequencies associated with the bosonic modes $\hat{a}$ and $\hat{b}$, respectively, and $\lambda(t)$ is the coupling strength between the two modes. The $\hat{a}$-mode is subject to dissipation, which is captured in the master equation by the dissipator $\mathcal{D}\hat{\rho}=\kappa\left(2\hat{a}\hat{\rho}\hat{a}^{\dagger}-\left\{\hat{a}^{\dagger}\hat{a},\hat{\rho}\right\}\right)$. The COS has an extensive application in multiple settings, such as but not limited to cavity-magnon systems [45, 46], atom-cavity systems [41, 42, 47, 48], and spin systems [49, 50].
The open COS has two phases: the NP which corresponds to a steady state with $a=b=0$, and an unbounded state where both modes exponentially diverge as $t\rightarrow\infty$ [41]. When nonlinearity is present, the unbounded state can be associated with a symmetry-broken phase, in which the modes choose a new steady state depending on their initial values. These two states are separated by the critical point [41]
@@
To show that the COS is a minimal model that can exhibit the KZM and its subsequent breakdown at small $\tau_{q}$, we consider its dynamics as it transitions from the NP to the unbounded state. We do this by initializing the system near the steady state of NP, $a_{0}=-b_{0}=0.01$. We then apply the quench protocol in Eq. (3) onto the COS and track the dynamics of the occupation number of the $\hat{a}$-mode, $|a|^{2}$. We finally determine $\hat{t}_{\mathrm{th}}$ by identifying the time it takes for $|a|^{2}$ to reach the threshold value, $|a|^{2}_{\mathrm{th}}$, after the ramp passes $\varepsilon(t=0)=0$. The crossover point at the transition time is then inferred back from $\hat{t}_{\mathrm{th}}$ using Eq. (3). We present in Fig. 2(b) the scaling of $\hat{t}_{\mathrm{th}}$ and $\varepsilon(\hat{t}_{\mathrm{th}})$ as a function of $\tau_{q}$. We can observe that for large $\tau_{q}$, or slow quench, all relevant quantities follow the power-law scaling predicted by the KZM. As we decrease $\tau_{q}$, $\varepsilon(\hat{t}_{\mathrm{th}})$ begins to saturate at a larger critical quench time, $\tau_{q,c}^{*}$, than $\hat{t}_{\mathrm{th}}$, as indicated by the solid line in Fig. 2(b). Finally, as $\tau_{q}\rightarrow 0$, $\hat{t}_{\mathrm{th}}$ approaches a constant value after passing another critical quench time, $\tau_{q,c}$, denoted in Fig. 2(b) as a dashed line. Note that the fluctuations in the scaling of $\hat{t}_{\mathrm{th}}$ and $\varepsilon(\hat{t}_{\mathrm{th}})$ can be attributed to the mean-field approach, which neglects any quantum fluctuation in the system’s dynamics. The scaling behaviour of $\hat{t}_{\mathrm{th}}$ and $\varepsilon(\hat{t}_{\mathrm{th}})$ implies that within the range $\tau_{q,c}<\tau_{q}\leq\tau_{q,c}^{*}$, there exists an intermediate regime between the true breakdown and the validity of the KZM, wherein the KZM remains valid even though the system appears to relax well after the quench has terminated. As shown in Fig. 2(c), this intermediate regime vanishes as $\kappa\rightarrow 0$, highlighting that this is a dissipation-induced effect.
@@
Figure 3: (a) (Top panel) Ramp protocol for $\omega\tau_{q}=500$ and $\varepsilon_{f}=0.2$. The circle point indicates $\varepsilon(\hat{t})$ while the diamond point marks $\varepsilon(\hat{t}_{\mathrm{th}})$. (Bottom panel) Exemplary dynamics of $|a|^{2}$ in the $\tau_{q}$-regime of the delayed KZM for $\kappa=0.1\omega$. The vertical dashed lines correspond to $\omega\hat{t}$, while the solid lines represent $\omega\hat{t}_{\mathrm{th}}$ for a threshold of $|a|^{2}_{\mathrm{th}}=2$. (b-c) potential surface of the open COS for (b) $\varepsilon(t)<0$, (c) $\varepsilon(t)=0$, and (d) $\varepsilon(t)>0$.
@@
We can understand this apparent contradiction between the scaling of the $\hat{t}_{\mathrm{th}}$ and $\varepsilon(\hat{t}_{\mathrm{th}})$ by looking at the dynamics of $|a|^{2}$ as the ramp crosses over $\varepsilon=0$. In Fig. 3(a), we present an exemplary dynamics of $|a|^{2}$ in the logarithmic scale for the regime $\tau_{q,c}<\tau_{q}\leq\tau_{q,c}^{*}$. Notice that before the system enters the unbounded state, $|a|^{2}$ first exponentially decays towards its steady state, indicating that the system does not freeze in the impulse regime. This dynamics is reminiscent of systems relaxing towards the global minimum of their energy surface due to dissipation, as sketched in Fig. 1(b). We can further establish this connection by obtaining the potential surface of the COS, which we can do by substituting the pseudo-position and momentum operators for the $\hat{a}$- mode
@@
is the effective potential of the COS in the closed system limit, $\kappa=0$. In this limit, the potential surface has a global minimum at $\hat{x}=\hat{y}=0$ when $\lambda<\lambda_{c}=\sqrt{\omega\omega_{0}}/2$, as shown in Fig. 3(b). It then loses its global minimum when $\lambda=\lambda_{c}$ as sketched in Fig. 3(c). Finally, the global minimum becomes a saddle point when $\lambda>\lambda_{c}$, as shown in Fig. 3(d). Note that in the presence of dissipation, the COS effective potential only becomes modified such that the critical point becomes Eq. 5, while the structure of the potential surface remains the same due to Eq. (8) being quadratic.
@@
With the above picture, we can now interpret the relaxation mechanism observed in Fig. 3(a) as follows. Suppose that we initialize our system such that $\lambda<\lambda_{c}$ and the initial states of $\hat{a}$ and $\hat{b}$ modes are close to the global minimum of $\hat{V}$. In the mean-field level, if $\kappa=0$, we can expect that the system will oscillate around the global minimum of $\hat{V}$ as we increase $\lambda$ using the finite ramp protocol defined in Eq. (3), together with the modification of the COS potential surface. As we cross $\lambda_{c}$, the global minimum of $\hat{V}$ becomes a saddle point. As such, any deviation of the initial state from the origin would eventually push the system to either the positive $\hat{x}$ and $-\hat{y}$ direction or vice versa, signalling the spontaneous symmetry breaking of the system.
@@
In the presence of dissipation, however, the system can still relax to the global minimum before the quench reaches $\lambda_{c}$ for sufficiently large quench time scales $\tau_{q}>\tau_{q,c}$. As a result, the slow deformation of the effective potential allows for the system to remain near $a=b=0$ even after passing the critical point where the potential loses its global minimum. The nudge from the system’s initial state eventually pushes the system towards a new minimum as the quench progresses, signalling the phase transition. This approach, however, only becomes detectable when $|a|^{2}$ reaches $|a|_{\mathrm{th}}^{2}$, which occurs only after the linear ramp has terminated. Thus, we observe the saturation of the crossover point at $\varepsilon_{f}$ even though $\hat{t}_{\mathrm{th}}$ follows the predicted scaling of the KZM, which hints that the system entered the adiabatic regime within the duration of the ramp. Note that the relaxation mechanism is not present in the closed limit, as hinted by the regime vanishing in Fig. 2(c) as $\kappa\rightarrow 0$. The delay between $\hat{t}$ and $\hat{t}_{\mathrm{th}}$ at finite $\tau_{q}$ motivates us to call this phenomenon delayed KZM.
@@
In the next section, we will show that the delayed KZM is a generic feature of open systems that can be mapped onto an effective COS. Moreover, we will demonstrate that not only is the delayed KZM induced purely by dissipation, but it also becomes more prominent when the system is quenched near criticality, $\varepsilon_{f}\approx 0$.

### v2

@@ Abstract
We report an intermediate regime in the quench time, $\tau_{q}$, separating the usual validity of the Kibble-Zurek mechanism (KZM) and its breakdown for rapid quenches in open systems under finite quench protocols. It manifests in the power-law scaling of the transition time with $\tau_{q}$ as the system appears to enter the adiabatic regime, even though the ramp is already terminated and the final quench value is held constant. This intermediate regime, which we dub the delayed KZM, emerges due to the dissipation, preventing the system from freezing in the impulse regime. This results in a large delay between the actual time the system undergoes a phase transition and the time inferred from a threshold-based criterion for the order parameter, as done in most experiments. We demonstrate using the open Dicke model and its one-dimensional lattice version that this phenomenon is a generic feature of open systems that can be mapped onto an effective coupled oscillator model. We also show that the phenomenon becomes more prominent near criticality, and its effects on the transition time measurement can be further exacerbated by large threshold values for an order parameter. Due to this, we propose an alternative method for threshold-based criterion which uses the spatiotemporal information, such as the system’s defect number, for identifying the transition time.
@@ I Introduction
Initially formulated to describe the evolution of topological defects in the early universe [1, 2, 3], the Kibble-Zurek Mechanism (KZM) has been successful in describing the dependence of the defect number and duration of a continuous phase transition on the quench timescale, $\tau_{q}$ [4]. In particular, the theory has been tested in multiple platforms, ranging from atomic Bose-Einstein condensates [5, 6, 7, 8, 9, 10, 11, 12, 13], spin systems [14, 15, 16, 17], Rydberg atom setups [18, 19], and trapped-ion systems [20, 21]. It has also been tested in dissipative quantum systems [22, 23, 24, 25, 26, 27, 17, 28], and has recently been extended to include generic nonequilibrium systems [29, 30, 31].
Under the standard KZM, a generic closed system with a continuous phase transition has a diverging relaxation time, $\tau$, and correlation length, $\xi$, as it approaches its critical point, $\lambda_{c}$. In particular, one expects $\tau$ and $\xi$ to scale as $\tau\propto|\varepsilon|^{-vz}$ and $\xi\propto|\varepsilon|^{-v}$ [4], respectively, where $\varepsilon=\left(\lambda-\lambda_{c}\right)/\lambda_{c}$ is the reduced distance of the control parameter, $\lambda$, from the critical point, while $v$ and $z$ are the static and dynamic critical exponents, respectively. It is then expected that if the system is linearly quenched via a ramp protocol, $\varepsilon=t/\tau_{q}$, the system will become frozen near $\lambda_{c}$ due to $\tau$ diverging. This motivates the introduction of the adiabatic-impulse (AI) approximation, where the system’s dynamics are classified into two regimes [4]. Far from $\lambda_{c}$, the system is in an adiabatic regime, in which its macroscopic quantities adiabatically follow the quench. Near $\lambda_{c}$, the system enters the impulse regime, wherein all relevant observables remain frozen even after passing $\lambda_{c}$. It only reenters the adiabatic regime and transitions to a new phase after some finite time referred to as the freeze-out time, $\hat{t}$, has passed [4]. This occurs after the system reaches the AI crossover point, $\varepsilon(\hat{t})$, setting $\hat{t}\sim\tau\left(\varepsilon(\hat{t})\right)$ [4]. The KZM predicts that, due to the scaling of $\tau$, $\hat{t}$ and $\varepsilon(\hat{t})$ must follow the scaling laws [4]
@@
Measuring the exact value of $\hat{t}$ and $\varepsilon(\hat{t})$ is a nontrivial task unlike defect counting due to the limitations in detecting the exact time a system reenters the adiabatic regime. As such, it is common to employ a threshold criterion and measure instead the transition time, $\hat{t}_{\mathrm{th}}$, which is the time it takes for an order parameter to reach a given threshold after passing $\lambda_{c}$. The crossover point at the transition time, $\varepsilon(\hat{t}_{\mathrm{th}})$, is similarly defined. For a sufficiently small threshold value, it is assumed that $\hat{t}_{\mathrm{th}}$ is a good approximation for $\hat{t}$. While this method is successful in showing the power-law scaling of $\hat{t}_{\mathrm{th}}$ and $\varepsilon(\hat{t}_{\mathrm{th}})$ as a function of $\tau_{q}$ [5, 12, 37, 10, 32, 28, 41], it remains unclear whether the inherent deviation between $\hat{t}$ and $\hat{t}_{\mathrm{th}}$ does not lead to any significant effects on the scaling of the KZM quantities for any generic quench protocols.
In this paper, we report an intermediate regime separating the breakdown and validity of the KZM appearing in open systems under a finite quench protocol depicted in Fig. 1(a). In this regime, the transition time follows the power-law scaling predicted by the KZM even though the system appears to relax after the quench has terminated, as illustrated in Fig. 1(b). As we will show later, this regime manifests precisely due to the dissipation exacerbating the deviation between the freeze-out time and the transition time, leading to a delay in the detection of the phase transition, as schematically represented in Fig. 1(b). We demonstrate using the open Dicke model (DM) [42, 43] and its one-dimensional lattice extension, the open Dicke lattice model (DLM) [44] that the range of $\tau_{q}$ where we observe this “delayed” KZM is a generic feature of open systems with finite dissipation strength, $\kappa$. We also show that the delayed KZM is more prominent near criticality and that its signatures become more significant for large threshold values for an order parameter. Thus, our paper highlights subtleties of the KZM in open systems in finite quench scenarios relevant to experiments.
@@ II Delayed KZM: Theory
where $\mathcal{D}\hat{\rho}=\sum_{\ell}\kappa_{\ell}\left(2\hat{L}_{\ell}\hat{\rho}\hat{L}_{\ell}^{\dagger}-\left\{\hat{L}^{\dagger}_{\ell}\hat{L}_{\ell},\hat{\rho}\right\}\right)$ is the dissipator and $\hat{H}(\varepsilon(t))$ is the time-dependent Hamiltonian of the system. The system undergoes a phase transition from a normal phase (NP), in which the global symmetry of the system is preserved, to a symmetry-broken phase via a finite quench protocol,
$\varepsilon(t)=\begin{cases}t/\tau_{q}\;\;\;&t_{i}\leq t\leq\varepsilon_{f}\tau_{q}\\ \varepsilon_{f}\;\;\;&\varepsilon_{f}\tau_{q}<t\leq t_{f}\end{cases}$ (3)
@@
where $\omega$ and $\omega_{0}$ are the transition frequencies associated with the bosonic modes $\hat{a}$ and $\hat{b}$, respectively, and $\lambda(t)$ is the coupling strength between the two modes. The $\hat{a}$ mode is subject to dissipation, which is captured in the master equation by the dissipator $\mathcal{D}\hat{\rho}=\kappa\left(2\hat{a}\hat{\rho}\hat{a}^{\dagger}-\left\{\hat{a}^{\dagger}\hat{a},\hat{\rho}\right\}\right)$. The COS has an extensive application in multiple settings, such as—but not limited—to cavity-magnon systems [46, 47], atom-cavity systems [42, 43, 48, 49], and spin systems [50, 51].
The open COS has two phases: the NP which corresponds to a steady state with $a=b=0$, and an unbounded state where both modes exponentially diverge as $t\rightarrow\infty$ [42]. When nonlinearity is present, the unbounded state can be associated with a symmetry-broken phase, in which the modes choose a new steady state depending on their initial values. These two states are separated by the critical point [42]:
@@
To show that the COS is a minimal model that can exhibit the KZM and its subsequent breakdown at small $\tau_{q}$, we consider its dynamics as it transitions from the NP to the unbounded state. We do this by initializing the system near the steady state of NP, $a_{0}=-b_{0}=0.01$. We then apply the quench protocol in Eq. (3) onto the COS and track the dynamics of the occupation number of the $\hat{a}$ mode, $|a|^{2}$. We finally determine $\hat{t}_{\mathrm{th}}$ by identifying the time it takes for $|a|^{2}$ to reach the threshold value, $|a|^{2}_{\mathrm{th}}$, after the ramp passes $\varepsilon(t=0)=0$. The crossover point at the transition time is then inferred back from $\hat{t}_{\mathrm{th}}$ using Eq. (3). We present in Fig. 2(b) the scaling of $\hat{t}_{\mathrm{th}}$ and $\varepsilon(\hat{t}_{\mathrm{th}})$ as a function of $\tau_{q}$. We can observe that for large $\tau_{q}$, or slow quench, all relevant quantities follow the power-law scaling predicted by the KZM. As we decrease $\tau_{q}$, $\varepsilon(\hat{t}_{\mathrm{th}})$ begins to saturate at a larger critical quench time, $\tau_{q,c}^{*}$, than $\hat{t}_{\mathrm{th}}$, as indicated by the solid line in Fig. 2(b). Finally, as $\tau_{q}\rightarrow 0$, $\hat{t}_{\mathrm{th}}$ approaches a constant value after passing another critical quench time, $\tau_{q,c}$, denoted in Fig. 2(b) as a dashed line. Note that the fluctuations in the scaling of $\hat{t}_{\mathrm{th}}$ and $\varepsilon(\hat{t}_{\mathrm{th}})$ can be attributed to the mean-field approach, which neglects any quantum fluctuation in the system’s dynamics. The scaling behavior of $\hat{t}_{\mathrm{th}}$ and $\varepsilon(\hat{t}_{\mathrm{th}})$ implies that within the range $\tau_{q,c}<\tau_{q}\leq\tau_{q,c}^{*}$, there exists an intermediate regime between the true breakdown and the validity of the KZM, wherein the KZM remains valid even though the system appears to relax well after the quench has terminated. As shown in Fig. 2(c), this intermediate regime vanishes as $\kappa\rightarrow 0$, highlighting that this is a dissipation-induced effect.
@@
Figure 3: (a) (top panel) Ramp protocol for $\omega\tau_{q}=500$ and $\varepsilon_{f}=0.2$. The circle point indicates $\varepsilon(\hat{t})$ while the diamond point marks $\varepsilon(\hat{t}_{\mathrm{th}})$. Bottom panel: Exemplary dynamics of $|a|^{2}$ in the $\tau_{q}$-regime of the delayed KZM for $\kappa=0.1\omega$. The vertical dashed lines correspond to $\omega\hat{t}$, while the solid lines represent $\omega\hat{t}_{\mathrm{th}}$ for a threshold of $|a|^{2}_{\mathrm{th}}=2$. (b), (c) potential surface of the open COS for (b) $\varepsilon(t)<0$, (c) $\varepsilon(t)=0$, and (d) $\varepsilon(t)>0$.
@@
We can understand this apparent contradiction between the scaling of the $\hat{t}_{\mathrm{th}}$ and $\varepsilon(\hat{t}_{\mathrm{th}})$ by looking at the dynamics of $|a|^{2}$ as the ramp crosses over $\varepsilon=0$. In Fig. 3(a), we present an exemplary dynamics of $|a|^{2}$ in the logarithmic scale for the regime $\tau_{q,c}<\tau_{q}\leq\tau_{q,c}^{*}$. Notice that before the system enters the unbounded state, $|a|^{2}$ first exponentially decays towards its steady state, indicating that the system does not freeze in the impulse regime. This dynamics is reminiscent of systems relaxing towards the global minimum of their energy surface due to dissipation, as sketched in Fig. 1(b). We can further establish this connection by obtaining the potential surface of the COS, which we can do by substituting the pseudoposition and momentum operators for the $\hat{a}$ mode,
@@
is the effective potential of the COS in the closed system limit, $\kappa=0$. In this limit, the potential surface has a global minimum at $\hat{x}=\hat{y}=0$ when $\lambda<\lambda_{c}=\sqrt{\omega\omega_{0}}/2$, as shown in Fig. 3(b). It then loses its global minimum when $\lambda=\lambda_{c}$ as sketched in Fig. 3(c). Finally, the global minimum becomes a saddle point when $\lambda>\lambda_{c}$, as shown in Fig. 3(d). Note that in the presence of dissipation, the COS effective potential only becomes modified such that the critical point becomes Eq. (5), while the structure of the potential surface remains the same due to Eq. (8) being quadratic.
@@
With the above picture, we can now interpret the relaxation mechanism observed in Fig. 3(a) as follows. Suppose that we initialize our system such that $\lambda<\lambda_{c}$ and the initial states of $\hat{a}$ and $\hat{b}$ modes are close to the global minimum of $\hat{V}$. In the mean-field level, if $\kappa=0$, we can expect that the system will oscillate around the global minimum of $\hat{V}$ as we increase $\lambda$ using the finite ramp protocol defined in Eq. (3), together with the modification of the COS potential surface. As we cross $\lambda_{c}$, the global minimum of $\hat{V}$ becomes a saddle point. As such, any deviation of the initial state from the origin would eventually push the system to either the positive $\hat{x}$ and $-\hat{y}$ direction or vice versa, signaling the spontaneous symmetry breaking of the system.
@@
In the presence of dissipation, however, the system can still relax to the global minimum before the quench reaches $\lambda_{c}$ for sufficiently large quench timescales $\tau_{q}>\tau_{q,c}$. As a result, the slow deformation of the effective potential allows for the system to remain near $a=b=0$ even after passing the critical point where the potential loses its global minimum. The nudge from the system’s initial state eventually pushes the system towards a new minimum as the quench progresses, signaling the phase transition. This approach, however, only becomes detectable when $|a|^{2}$ reaches $|a|_{\mathrm{th}}^{2}$, which occurs only after the linear ramp has terminated. Thus, we observe the saturation of the crossover point at $\varepsilon_{f}$ even though $\hat{t}_{\mathrm{th}}$ follows the predicted scaling of the KZM, which hints that the system entered the adiabatic regime within the duration of the ramp. Note that the relaxation mechanism is not present in the closed limit, as hinted by the regime vanishing in Fig. 2(c) as $\kappa\rightarrow 0$. The delay between $\hat{t}$ and $\hat{t}_{\mathrm{th}}$ at finite $\tau_{q}$ motivates us to call this phenomenon delayed KZM.
@@
In the next section, we will show that the delayed KZM is a generic feature of open systems that can be mapped onto an effective COS. Moreover, we will demonstrate that not only is the delayed KZM induced purely by dissipation but it also becomes more prominent when the system is quenched near criticality, $\varepsilon_{f}\approx 0$.